OpenStudy (anonymous):

find three consecutive even integers so that the twice the sum of the second and third is twelve less that six times the first a. 9, 10, 11 b. 6, 8, 10 c. 2, 4, 6 d. 3, 4, 5 e. none of the above

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

3 consecutive even integers (consecutive evens are split by factors of 2) x, x+2, x+4 Automatically we can cross out a and d

OpenStudy (bibby):

now translate the words into an equation twice the sum of the second and third is twelve less that six times the first

OpenStudy (anonymous):

that's the part i was having trouble with, my mind is not grasping translating this problem. which is basically, the entirety of the problem.

OpenStudy (alekos):

let the three integers be x,y and z so y=x+1 and z=x+2

OpenStudy (bibby):

ok, well I'll help you understand the idea before moving on. Consecutive means one after another. Even means divisible by 2. so 2,4,6 or 6,8,10 or 4,6,8 now we want to represent this relationship using a variable x = 2 then y and z would be 4 and 6 or x+2 and x+4 respectively

OpenStudy (alekos):

so according to the problem 2(y+z) = 6x -12

OpenStudy (bibby):

@alekos "consecutive even integers"

OpenStudy (alekos):

that's fine. then y=x+2 and z=x+4 so just substitute y & z into the equation and solve for x

OpenStudy (anonymous):

alright, that's starting to make a little more sense. this problem i have no idea even where to begin so you explaining it is helping a ton

OpenStudy (bibby):

well the equation has been written out soooo

OpenStudy (anonymous):

i know i'm solving it now

OpenStudy (bibby):

oh yeah that wasn't meant to sound impatient

OpenStudy (anonymous):

i'm getting B. 6, 8, 10

OpenStudy (alekos):

that's not what I got

OpenStudy (bibby):

2(y+z) = 6x -12 y=x+2 z=x+4 2(x+2+x+4)=6x-12 2(x+6)=6x-12 2x+12=6x-12 subtract 2x and add 12 24=4x x=6

OpenStudy (alekos):

5th line should be 2(2x+6) = 6x-12

OpenStudy (bibby):

yeah it should be 12,14,16 my bad